OpenStudy (anonymous):

order from widest to narrowest, y=-3x^2, y=-4x^2, y=-5x^2? HELP

OpenStudy (anonymous):

the order is as it is listed... the widest is the first the thinnest is the last.

OpenStudy (anonymous):

the bigger the coefficient of the x^2 term, the thinner it gets.

OpenStudy (anonymous):

** the bigger the absolute value....

OpenStudy (campbell_st):

the larger the absolute value of the coefficient of x^2 the narrower the gap... in order the narrowest is y = -5x^2 |-5| = 5 y = -4x^2 |-4|= 4 etc so y = -1/2x^2 is a wide parabola